OpenStudy (anonymous):

Find the area of a parallelogram with base b and height h. b = 82 cm h = 16.6 cm A)1361.2 cm2 B)6,724 cm2 C)137.78 cm2 D)98.6 cm2

OpenStudy (merchandize):

area of parallelogram=base*height so multiply base and height and you will get the area of parallelogram.

OpenStudy (merchandize):

i.e 82*16.6=161.2cm^2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what is it

jimthompson5910 (jim_thompson5910):

82*16.6 = 1361.2 and it's not 82*16.6=161.2

OpenStudy (merchandize):

sorry i just missed that 3 in between.. so it's 1361.2